Can you assist me with this answer: During periods of hyperinflation, prices increase rapidly over short periods of time. One of the worst cases of hyperinflation in history occurred in Zimbabwe in 2008 where prices doubled every day. If that rate continued for just 7 days, how much would a French baguette cost after 7 days if the cost at the beginning of the week was $0.6? The cost after 7 days will be
Answer #1

Inflation rate=100% (i.e. prices doubled) per day

Initial price=Po=$0.6

Number of periods=n=7 days

Price after 7 days=Po*(1+g)7=0.60*(1+100%)7=$76.80